OpenStudy (anonymous):

Find another name for given number, but with denominator indicated. Use multiplication by 1. 3/8=?/32

OpenStudy (lncognlto):

\[8 \times x = 32\]Therefore\[x = 4\]Now\[3 \times 4 = ?\]

OpenStudy (anonymous):

12/32

OpenStudy (lastdaywork):

1. 3/8=?/32 Does . implies multiplication or decimal??